OpenStudy (anonymous):

Integration @ganeshie8

OpenStudy (anonymous):

\[Given~\int\limits\limits_{-1}^{2}[f(x)+1]~dx=9~and~\int\limits\limits_{2}^{5}f(x)~dx=16,find\]

OpenStudy (anonymous):

\[a)\int\limits_{-1}^{2}3f(x)~dx\]

OpenStudy (anonymous):

\[b)\int\limits_{-1}^{5}[2f(x)-1]dx\]

ganeshie8 (ganeshie8):

integral has below properties : \[\int \color{red}{k}f(x) \, dx = \color{red}{k}\int f(x)\,dx\] \[\int [f(x)+g(x)] \, dx = \int f(x)\,dx+\int g(x)\,dx\]

ganeshie8 (ganeshie8):

We're given : \[\int\limits\limits_{-1}^{2}[f(x)+1]~dx=9\] using the previously listed integral properties, can you find the value of \(\int\limits_{-1}^2f(x)\, dx\) ?

OpenStudy (anonymous):

\[\int\limits\limits_{-1}^{2}f(x)+\int\limits_{-1}^{2}1~dx\]

ganeshie8 (ganeshie8):

Yes, keep going...

ganeshie8 (ganeshie8):

don't forget the dx in first integral

ganeshie8 (ganeshie8):

\[\int\limits\limits_{-1}^{2}f(x)\,\color{red}{dx}+\int\limits_{-1}^{2}1~dx\]

OpenStudy (anonymous):

\[[\frac{ x^2 }{ 2 }]_{-1}^{2}+[x]_{-1}^{2}=9\]

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

we don't know what f(x) is so we cannot integrate it directly

OpenStudy (anonymous):

\[\int\limits_{-1}^{2}f(x)~dx=6\]

ganeshie8 (ganeshie8):

We're given : \(\int\limits\limits_{-1}^{2}[f(x)+1]~dx=9\) \(\int\limits\limits_{-1}^{2}f(x)~dx+\int\limits\limits_{-1}^{2}1~dx=9\) \(\int\limits\limits_{-1}^{2}f(x)~dx = 9 - \int\limits\limits_{-1}^{2}1~dx\) \(\int\limits\limits_{-1}^{2}f(x)~dx = 9 - 3 = 6\)

ganeshie8 (ganeshie8):

So, we have two things : \[\int\limits_{-1}^{2}f(x)~dx=6\tag{1}\] \[\int\limits_{2}^{5}f(x)~dx=16\tag{2}\] we're ready to answer the questions

ganeshie8 (ganeshie8):

\[a)\int\limits_{-1}^{2}3f(x)~dx\] do you want to try working this ? you just need to use one of the previously listed integral properties

OpenStudy (anonymous):

okay,i will try

OpenStudy (anonymous):

\[3\int\limits_{-1}^{2}f(x)~dx=3(6)=18\]

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

for question b)\[\int\limits_{-1}^{5}[2x-1]~dx=2\int\limits_{-1}^{5}[f(x)-\frac{ 1 }{ 2 }]dx\] is it correct?

OpenStudy (anonymous):

\[\int\limits_{-1}^{5}f(x)=22\]

ganeshie8 (ganeshie8):

that is correct, but i feel splitting like below looks better : \[\int\limits_{-1}^{5}[2f(x)-1]~dx=\int\limits_{-1}^{5}2f(x)\,dx-\int\limits_{-1}^{5}1\,dx\\~\\~\\ =2\int\limits_{-1}^{5}f(x)\,dx-\int\limits_{-1}^{5}1\,dx\\~\\~\\\]

ganeshie8 (ganeshie8):

\[\int\limits_{-1}^{5}f(x)=22\] how did u get this ?

OpenStudy (anonymous):

i think i did wrong :\ 6+16=22

ganeshie8 (ganeshie8):

that is correct! i just wanted to know how you worked it, thats all :)

OpenStudy (anonymous):

oh i c... :D

ganeshie8 (ganeshie8):

As you can see, definite integrals just add up like areas : \[\int\limits_a^{\color{red}{c}}f(x)\,dx + \int\limits_{\color{red}{c}}^b f(x)\, dx = \int\limits_a^bf(x)\, dx\]

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

(Area under curve between a to c ) + (Area under curve between c to b ) = (Area under curve between a to b )

OpenStudy (anonymous):

\[\int\limits_{-1}^{5}[2f(x)-1]=44-6=38\]

OpenStudy (anonymous):

Thank you @ganeshie8 :)

ganeshie8 (ganeshie8):

Looks good! remember, this is calculus, knowing how to work some particular problem is not enough... you need to have very good understanding of graphs of functions to really enjoy calculus

OpenStudy (anonymous):

okay, thnx @ganeshie8